what is the a,b,and c for y= -x^2+2 and the x intercepts?
Answer by ewatrrr(24785)

 
Hi
using standard form: y = ax^2 + bx + c
y= -x^2+2 || a = -1 , b = 0 and c = 2
x -intercepts (when y = 0)
-x^2 + 2 = 0
x^2 = 2
x = ± sqrt(2)